OpenStudy (hiyathere_20):

The slope of the line passing through the points (0, 10) and (0, 3) is 13. TRUE OR FALSE

jabez177 (jabez177):

\[\frac{ y^2-y^1 }{ x^2-x^1 }\]

OpenStudy (mathmale):

You are given 2 poins. Use the slope formula to calculate the actual slope, and then compare your result to the given m=13. Unfortunately, \[\frac{ y^2-y^1 }{ x^2-x^1 }\]

jabez177 (jabez177):

Then we just plug and chug. Plug in the numbers where the variables are and solve. :)

OpenStudy (hiyathere_20):

tRUE ?

OpenStudy (mathmale):

is the wrong formula; those 1's and 2's indicate exponentiation, which is not correct. Rather, please use\[m=\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]

OpenStudy (hiyathere_20):

I got -7? so it is false correct??????

OpenStudy (mathmale):

Hiya, hiya: please share your work.

OpenStudy (skullpatrol):

Note that when both x-coordinates are equal the slope is undefined.

OpenStudy (hiyathere_20):

I just subtracted 3-10 which gave me -7

OpenStudy (mathmale):

Look carefully at the problem statement: "the points (0, 10) and (0, 3)" shows that x does not change, but y changes. In such situations, what could be said of the slope? Read skullpatrol's comment, above.

OpenStudy (hiyathere_20):

OKAY SO FALSE????

OpenStudy (skullpatrol):

Yes, false.

OpenStudy (mathmale):

-7 for what? Please, reread the original question. For what does it ask you?

OpenStudy (hiyathere_20):

thankkkkkuuuuuuu SKULLPATROL AND EVERYONE ELSE FOR THE HELP on to the next question :D

OpenStudy (mathmale):

\[m=\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]

OpenStudy (mathmale):

Post each question separately, please.

OpenStudy (skullpatrol):

Thanks for trying to learn :-)

OpenStudy (hiyathere_20):

I am Mathmale thanks!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (skullpatrol):

$$\huge \text{slope = } \dfrac{y_2 - y_1}{x_2 - x_1} \text{, } x_2 \ne x_1$$